# Towards Trapped-Ion Thermometry Using Cavity-Based EIT

### v2

@@ Abstract
We present a technique for measuring ion temperature using cavity-based electromagnetically induced transparency (EIT) applicable for cavity-qed systems in the strong coupling regime. This method enables efficient extraction of the ion’s phonon occupation number following sub-Doppler cooling close to the motional ground state. The proposed method relies on monitoring the cavity probe transmission while scanning the probe laser frequency once cavity EIT is established using the control beam, significantly simplifying the measurement procedure. We theoretically establish a model that demonstrates influence of thermal state of the trapped ion vis-a-vis the EIT linewidth measured. We show how the cavity EIT transmission may be used as a thermometry tool to deduce the ion temperature as well as the motional state for an ion in the sub-Doppler cooling regime even for systems that are in the weak coupling regime. The current method can only be used for operation in the resolved-sideband regime, where individual motional states can be selectively addressed for all relevant transitions either by selecting appropriate energy levels for the three-level system or by employing strong confinement with high secular frequencies ($\sim 10MHz$).
@@ I Introduction
Achieving and confirming near-ground-state cooling is essential, since even a small residual motional excitation can degrade coherence and reduce the fidelity of quantum operations [5, 6]. With successful demonstrations across a broad range of ion species and trap architectures, resolved sideband cooling has emerged as the standard technique for preparing trapped ions close to the motional ground state [7, 8, 5]. For this reason, accurate thermometry after sideband cooling is crucial for describing cooling performance and validating subsequent quantum operations. After compensation of excess micromotion, it is possible to characterize the motion of a trapped ion in a radio-frequency Paul trap [5]. Ion temperature is directly determined by a discrete phonon occupation number that characterizes the motional state. Determining this phonon occupation provides a crucial benchmark for coherent quantum operations and enables a direct evaluation of the ions’ thermal state [1, 5, 9].
@@
There has been a growing interest in cavity quantum electrodynamics with trapped ions due to its potential to realize scalable quantum networks, enhanced light-matter coupling, and effective ion-photon interfaces [13, 14, 15, 16, 17, 18, 19]. Interference-based phenomena such as cavity-assisted electromagnetic-induced transparency (EIT) are made possible by the presence of an optical cavity, which alters the interaction between the ion and the electromagnetic field. Since its discovery, EIT [20] has made it possible for a wide range of coherent optical control applications, such as slow light in atomic medium [21], quantum memory [22], laser cooling [23], precision spectroscopy, and sensing [13].
@@
Cavity-based EIT has been studied in a number of regimes to enable features such as narrowing of the cavity linewidth and enhanced optical nonlinearities [16, 17, 24]. A key feature of cavity-induced EIT is its sensitivity to decoherence mechanisms in the coupled ion-cavity system. Coupling to vibrational sidebands causes phonon-dependent dephasing of the effective $\Lambda$-type system in the presence of motional excitation, changing the coherence and causing the EIT interference [25, 23]. The cavity-EIT transparency window is thus systematically expanded by thermal phonons. Thermally induced fluctuations alter the effective decoherence rates in the presence of motional excitation which affects the EIT transparency features’ width and shape. This dependence suggests that the cavity-EIT can serve as an indirect probe of the ion’s motional temperature.
@@
In this paper, we theoretically investigate a cavity-enhanced thermometry scheme for trapped ions that exploits this phonon-induced modification of cavity-induced EIT. By including vibrational sidebands and the ion’s motion coupled to a thermal reservoir in a $\Lambda$-type ion-cavity setup, we demonstrate that the phonon occupation of the trapped ion systematically affects the linewidth of the cavity-EIT resonance. This dependence can be used as a direct probe of the mean phonon number by tuning the coupling field to motional sideband transitions. Therefore, the suggested method offers a cavity-compatible thermometry technique that does not depend on projective state detection after sideband cooling. In this theoretical framework a direct link between measurable modification of the cavity-EIT resonance and motional decoherence induced by thermal phonon is established. This enhances current thermometry methods and offers an experimentally feasible path for temperature diagnostics in trapped-ion cavity-QED systems.
@@
We show that the coupling laser causes a phonon number-dependent change in the effective Rabi frequency when it is tuned to motional sideband transition. This results in a systematic and temperature-dependent broadening of the cavity-EIT feature. We determine a quantitative relationship between the measured cavity-EIT linewidth and the ion temperature by numerically solving the complete open-system dynamics, which includes cavity loss, spontaneous emission, and phonon damping caused by a thermal reservoir. Our results indicate that this technique allows for precise thermometry in the resolved-sideband and sub-Doppler regimes and can be expanded to parameter regimes that are consistent with experimentally realized strong coupling ion-cavity systems[16, 17, 18, 19].
@@
In this work a direct link between motional decoherence and cavity transmission properties is established, and a cavity-compatible minimally invasive thermometry method for trapped ions is presented. Beyond thermometry, the shown sensitivity of cavity-EIT to the motional state of the ion creates new opportunities for studying measurement-induced back-action in cavity-QED systems and for ongoing heating dynamics monitoring.
The paper is organized as follows: In section II, we present the theoretical model and Hamiltonian including the treatment of ion motion and dissipative processes. In section III, numerical results for linewidth extraction and cavity transmission are shown. The section IV examines how phonons affect the cavity-EIT spectrum. The analysis of ion temperature and the role of thermalization dynamics are discussed in section V. Finally, a summary of the key results and concluding remarks are given in section section VIII.
@@ II Theory and model
Cavity-EIT systems have been extensively studied in the literature [16, 17, 26, 24, 27, 28, 29, 30]. These studies demonstrate that the cavity output can be effectively controlled by several key parameters, including the control field intensity ($\Omega_{c}$), the atom–cavity coupling strength ($g$), and the cavity decay rate ($\kappa$). Among these parameters, the control field intensity, $\Omega_{c}$, represents an experimentally accessible degree of freedom, whereas the others are largely determined by the cavity geometry, mirror reflectivity, and the choice of atomic energy levels. As a result, parameters such as $g$ and $\kappa$ cannot be easily tuned under typical experimental constraints, whereas $\Omega_{c}$ can be easily modified by tweaking the laser power. As the control field intensity $\Omega_{c}$ increases, the full width at half maximum (FWHM) of the EIT resonance also increases significantly, typically by approximately $10~\mathrm{kHz}$. This is the key on which the ion thermometry method is further established.
For the purpose of ion thermometry, it is essential to include the vibrational levels associated with each relevant electronic energy level. Addressing these vibrational states requires operation in the resolved-sideband regime ($\gamma\ll\omega_{sec}$) [31, 1], where the splitting between adjacent vibrational levels, $\hbar\omega_{sec}$, exceeds the natural linewidth of the corresponding transition, $2\gamma$. Here, $\gamma$ denotes the spontaneous decay rate of the transition, and $\omega_{sec}$ represents the secular frequency of the trapped ion.
Henceforth, a transition between two energy levels refers to a transition between their corresponding vibrational states. The allowed transitions can be classified into three main types: the carrier transition, the blue sideband (BSB) transition, and the red sideband (RSB) transition, corresponding to $\Delta n=0$ and $\Delta n=\pm 1$, respectively. The vibrational level occupancy, $n$, directly influences the ion–field coupling strength when sideband transitions are addressed. Consequently, the Rabi frequencies for the RSB and BSB transitions are modified as [1],
@@
In contrast, for the carrier transition, the Rabi frequency remains unchanged ($\Omega_{0}$). When the control laser is tuned to a sideband transition (BSB or RSB), the effect of the phonon occupancy, $n$, becomes evident in the EIT linewidth through its influence on the Rabi frequency, $\Omega_{c}$. Conversely, variations in the EIT linewidth provide an estimate of changes in the ion’s vibrational occupancy. Motivated by this relationship, we investigate this phenomenon through numerical simulations and identify a systematic approach to estimate the ion temperature, which is directly linked to the phonon occupancy.
@@
In a frame rotating at the probe frequency $\omega_{probe}$, and under the electric dipole and rotating-wave approximations (RWA), the time-independent Hamiltonian for the atom-cavity system is given by (with $\hbar$ =1):
@@
$\frac{d\rho}{dt}=-i\left[H,\rho\right]+\mathcal{L}_{\mathrm{cav}}[\rho]+\mathcal{L}_{\mathrm{atom}}[\rho]+\mathcal{L}_{\mathrm{ph}}[\rho],$ (3)
with cavity loss
$\mathcal{L}_{\mathrm{cav}}[\rho]=\kappa\left(2a\rho a^{\dagger}-a^{\dagger}a\rho-\rho a^{\dagger}a\right),$ (4)
where $\kappa$ denotes the decay rate of the cavity field.
@@
$\mathcal{L}_{\mathrm{atom}}[\rho]=\sum_{j=1,2}\gamma_{3j}\left(2\sigma_{j3}\rho\sigma_{3j}-\sigma_{33}\rho-\rho\sigma_{33}\right),$ (5)
[…]
The phonon states are inherent to damping due to fluctuating electric field noise present in real ion trap. To account this scenerio we considered the trapped ion system such that it is coupled to a thermal bath. The fluctuation causes, initially the phonon numbers to be exchanged between the ion and reservoir, attending an equilibrium condition. The damping of the vibrational mode of the ion due to its coupling with a thermal reservoir is described by [6]
$\mathcal{L}_{\mathrm{ph}}[\rho]=$ $\gamma_{b}(n_{\mathrm{th}}+1)\left(2b\rho b^{\dagger}-b^{\dagger}b\rho-\rho b^{\dagger}b\right)$ (6)
$+\gamma_{b}n_{\mathrm{th}}\left(2b^{\dagger}\rho b-bb^{\dagger}\rho-\rho bb^{\dagger}\right),$
where $\gamma_{b}$ is the phonon damping rate and $n_{\mathrm{th}}$ denotes the mean thermal phonon number at equilibrium. The average phonon number ($\bar{n}$) is related to temperature of the ion by,
@@
Following Equation 6 the rate equation for phonon can be derived as [6],
@@
So, close to ground state ($\bar{n}=0$), $\dot{\bar{n}}(t)=\gamma_{b}n_{th}$. This gives an initial estimation of heating rate of the ion. Equation 8, also shows that if a long time interval is taken, i.e. at equilibrium, $\bar{n}_{eq}=n_{th}$. So, determining the value of $n_{th}$ allows us to estimate the ion’s temperature owing to Equation 7.
@@ III Analysis of Transmission in Cavity-EIT
We numerically solve the master equation (Equation 3) using the Hamiltonian given in Equation 2. The probe-field detuning is varied across the resonance, and the steady-state density matrix is obtained for each detuning value. From the steady-state solution, the photon buildup inside the cavity is extracted by evaluating the expectation value of the photon number operator, $\langle a^{\dagger}a\rangle$. Multiplying the intra-cavity photon number by the cavity decay rate, $\kappa$, yields the cavity output, or the cavity-EIT spectrum ( see Figure 2). This spectrum exhibits the characteristic features of previously reported cavity-EIT systems [16, 33, 17, 24].
@@
The cavity linewidth is observed to be narrower than that of an empty cavity, and the vacuum Rabi splitting peaks appear at $\Delta=\pm\sqrt{g^{2}+\Omega_{c}^{2}}$. For all simulations presented in this work, unless otherwise stated, the system parameters are fixed as follows. All quantities are normalized with respect to the cavity decay rate, $\kappa$, which is taken to be $0.4~\mathrm{MHz}$, a typical value for a high-finesse cavity. The coupling-field Rabi frequency, $\Omega_{c}$, is set to $2.5\kappa$, while the atom–cavity coupling strength, $g$, is chosen as $3\kappa$. The decay rates from the excited states, $\gamma_{ge}$ and $\gamma_{ue}$, are taken to be equal to $\kappa$. In addition, motional-state decoherence is included to account for unwanted heating or cooling of the ion arising from phase noise and radiative fluctuations present in realistic ion traps. To model this effect, a weak coupling between the motional state and the environment is introduced, characterized by a coupling constant $\gamma_{b}\approx 0.25\kappa$.
@@ IV Effect of thermal state on cavity EIT linewidth
The usual cavity-EIT formalism discussed in the literature does not explicitly include the ion’s motional state. This inherently makes an assumption that the ion is fixed in space and the effect of ion thermalization doesn’t affect the cavity-EIT output. When the ion’s motion is not considered for cavity based EIT the Hamiltonian is simplified to,
@@
Where the different terms have their usual meaning as described in section II. We refer this formalism as non-thermal and the Hamiltonian given in Equation 9, the non-thermal cavity-EIT Hamiltonian. The output of the cavity is traced from the average of photon number operator with the variation of the probe field detuning. The same cavity transmission can also be analytically calculated from the expression[34],
@@
When the motional states are incorporated into the Hamiltonian (Equation 2), providing a more general description of a trapped ion, the cavity-EIT linewidth deviates from the non-thermal case. As the temperature increases, the thermal nature of the ion leads to a higher motional occupancy. This increase in motional occupation modifies the cavity-EIT linewidth by changing $\Omega_{c}$, as discussed in detail in section II.
Figure 4: (a) Cavity-EIT spectrum for different $n_{th}$ value (b) Ion temperature and average phonon occupancy ($\bar{n}$) as a function of cavity-EIT linewidth. Parameters used are: $\kappa=0.4MHz$, $g=3\kappa$, $\gamma_{eg}=\gamma_{eu}=\kappa$, $n_{th}=\langle n\rangle=1$, $\gamma_{b}=0.25\kappa$.
To compare the linewidths in the thermal and non-thermal cases, we perform simulations for both scenarios using the same values of $\Omega_{c}$ as those employed in the analytical plot in 3(a). The thermal simulations are carried out over a range of temperatures to capture the dependence of the cavity linewidth on the ion temperature. In 3(b), the thermal and non-thermal linewidths are presented, with the analytical linewidths from 3(a) used as the reference on the x-axis.
We observe that, over the entire range, the non-thermal linewidth remains lower than the thermal linewidth, and that the linewidth curves shift toward higher values as the temperature increases. This behavior provides further motivation for quantifying the ion temperature, as it modifies the cavity linewidth, which is highly sensitive to temperature-induced changes in the ion’s motional state.

### v3

@@ Abstract
We present a technique for measuring ion temperature using cavity-based electromagnetically induced transparency (EIT) applicable for cavity quantum electrodynamics (QED) systems in the strong coupling regime. This method enables efficient extraction of the ion’s phonon occupation number following sub-Doppler cooling close to the motional ground state. The proposed method relies on monitoring the cavity probe transmission while scanning the probe laser frequency to establish cavity-induced EIT using the control beam, significantly simplifying the measurement procedure. We establish a theoretical model that demonstrates the influence of the thermal state of the trapped ion vis-à-vis the EIT linewidth measured. We show through numerical simulations how the cavity-induced EIT transmission may be used as a thermometry tool to deduce the ion temperature as well as the motional state of the ion in the sub-Doppler cooling regime, even for systems that are in the weak coupling regime. The current method can only be used for operation in the resolved-sideband regime, where individual motional states can be selectively addressed for all relevant transitions either by selecting appropriate energy levels for the three-level system or by employing strong confinement with high secular frequencies ($\sim 10MHz$).
@@ I Introduction
With successful demonstrations across a broad range of ion species and trap architectures, resolved sideband cooling has emerged as the standard technique for preparing trapped ions close to the motional ground state [5, 6, 7]. Achieving and confirming near-ground-state cooling is essential, since even a small residual motional excitation can degrade coherence and reduce the fidelity of quantum operations [7, 8]. For this reason, accurate thermometry after sideband cooling is crucial for describing cooling performance and validating subsequent quantum operations. After compensation of excess micromotion, it is possible to characterize the motion of a trapped ion in a radio-frequency Paul trap [7]. Ion temperature is directly determined by a discrete phonon occupation number that characterizes the motional state. Determining this phonon occupation provides a crucial benchmark for coherent quantum operations and enables a direct evaluation of the ions’ thermal state [1, 7, 9].
@@
Over the past few decades, following the seminal works of Kimble [13], Haroche [14] and Rempe [15], there has been growing interest in cavity quantum electrodynamics with trapped ions due to its potential for the realization of scalable quantum networks, enhanced light-matter coupling, and effective ion-photon interfaces for quantum information processing schemes and applications [16, 17, 18, 19, 20, 21, 22]. Interference-based phenomena such as cavity-assisted electromagnetic-induced transparency (EIT) are made possible by the presence of an optical cavity, which alters the interaction between the ion and the electromagnetic field. Since its discovery, EIT [23] has made it possible for a wide range of coherent optical control applications, such as slow light in atomic medium [24], quantum memory [25], laser cooling [26], precision spectroscopy, and sensing [16].
@@
Cavity-based EIT has been studied in a number of regimes to enable features such as narrowing of the cavity linewidth and enhanced optical nonlinearities [19, 20, 27]. A key feature of cavity-induced EIT is its sensitivity to decoherence mechanisms in the coupled ion-cavity system. Coupling to vibrational sidebands causes phonon-dependent dephasing of the effective $\Lambda$-type system in the presence of motional excitation, changing the coherence, and altering the EIT interference [28, 26]. The cavity-EIT transparency window is thus systematically expanded by thermal phonons. Thermally induced fluctuations alter the effective decoherence rates in the presence of motional excitation which affects the EIT transparency features’ width and shape. This dependence suggests that the cavity-EIT can serve as an indirect probe of the ion’s motional temperature.
@@
In this article, we theoretically investigate a cavity-enhanced thermometry scheme for trapped ions that exploits the phonon-induced modification of cavity-induced EIT. We consider a $\Lambda$-type trapped ion cavity-QED system in which an auxiliary internal transition of the ion is driven by a coupling (control) laser and the cavity mode is interacted with a weak probe field. By including vibrational sidebands and the ion’s motion coupled to a thermal reservoir in a $\Lambda$-type ion-cavity setup, we demonstrate that the phonon occupation of the trapped ion modifies the cavity EIT linewidth in a systematic manner. This dependence can be used as a direct probe of the mean phonon number by tuning the coupling field to motional sideband transitions. Thus, the suggested method offers a cavity-compatible thermometry technique that does not depend on projective state detection after sideband cooling. Our theoretical framework establishes a direct link between measurable modification of the cavity-EIT resonance and motional decoherence induced by thermal phonon. The proposed approach provides an experimentally feasible alternative to conventional sideband thermometry in trapped-ion cavity-QED systems.
@@
We show that the coupling laser causes a phonon number-dependent change in the effective Rabi frequency when it is tuned to a motional sideband transition. This results in a systematic and temperature-dependent broadening of the cavity-EIT feature. We determine a quantitative relationship between the measured cavity-EIT linewidth and the ion temperature by numerically solving the complete open-system dynamics, which includes cavity loss, spontaneous emission, and phonon damping caused by a thermal reservoir. Our results indicate that this technique allows for precise thermometry in the resolved-sideband and sub-Doppler regimes and can be expanded to parameter regimes that are consistent with experimentally realized strong coupling ion-cavity systems[19, 20, 21, 22].
We subsequently extend the model to the multi-ion system by modifying the atom-cavity interaction Hamiltonian to include a summation over all ions coupled to the same cavity mode. In this formulation, each ion interacts with the cavity field identically, and the collective response of the ion ensemble naturally emerges from the full many-body master equation. This approach allows us to examine thermometry in regimes where the cavity response is improved by the collective light-matter interaction and the single-ion coupling strength to the intracavity field places the system in the weak coupling regime.
@@
To summarize, we establish a direct link between motional decoherence and cavity transmission, and propose a cavity-compatible minimally invasive thermometry technique for trapped ions through modeling and simulations. Beyond thermometry, the demonstrated sensitivity of cavity-EIT to the motional state of the ion creates new opportunities for studying measurement-induced back-action in cavity-QED systems and for monitoring of thermal dynamics.
The article is organized as follows: In section II, we present the theoretical model and Hamiltonian, including the treatment of ion motion and dissipative processes. In section III, numerical results for linewidth extraction and cavity transmission are shown. The section IV examines how phonons affect the cavity-EIT spectrum. The analysis of ion temperature and the role of thermalization dynamics are discussed in section V. The discussion then extends to multi-ion cavity-QED systems operating in the weak single-ion coupling regime in section VI, where we demonstrate that thermometry is made possible even in cases where individual coupling strengths are small through collective enhancement. section VII further investigates the applicability of the proposed method to systems with large excited state decay rates. We show conclusively that thermometry remains feasible under realistic experimental conditions. We present a realistic experimental realization roadmap in section VIII covering parameter regimes compatible with existing ion-cavity platform trap requirements and viable cavity configurations. Finally, a summary of the key results and concluding remarks are given in section section IX.
@@ II Theory and model
Cavity-based EIT systems have been extensively studied over the past two decades [19, 20, 29, 27, 30, 31, 32, 33]. These studies demonstrate that the cavity output can be effectively controlled by several key parameters, including the atom-cavity coupling strength ($g$), control field Rabi frequency ($\Omega_{c}$), and the cavity decay rate ($\kappa$). Among these parameters, the control field intensity, $\Omega_{c}$, represents an experimentally accessible degree of freedom, whereas the others are primarily defined by the cavity geometry, mirror reflectivity, and the choice of atomic energy levels. As a result, parameters such as $g$ and $\kappa$ cannot be easily tuned under typical experimental constraints, whereas $\Omega_{c}$ can be easily modified by tweaking the laser power. As the control field intensity $\Omega_{c}$ increases, the full width at half maximum (FWHM) of the EIT resonance also increases significantly, typically by $\approx 10~\mathrm{kHz}$. This dependence forms the basis of the proposed thermometry scheme.
For the purpose of ion thermometry, it is essential to include the vibrational levels associated with each relevant electronic energy level. Addressing these vibrational states requires operation in the resolved-sideband regime ($\gamma\ll\omega_{sec}$) [34, 1], where the splitting between adjacent vibrational levels, $\hbar\omega_{sec}$, exceeds the natural linewidth of the corresponding transition, $2\gamma$. Here, $\gamma$ denotes the spontaneous decay rate of the transition and $\omega_{sec}$ represents the secular frequency of the trapped ion.
Henceforth, a transition between two energy levels refers to a transition between their corresponding vibrational states. The allowed transitions can be classified into three main types: the carrier transition, the blue sideband (BSB) transition, and the red sideband (RSB) transition, corresponding to $\Delta n=0$ and $\Delta n=\pm 1$, respectively. When sideband transitions are addressed, the vibrational level occupancy, n, certainly influences the ion-field coupling strength. Consequently, the Rabi frequencies for the RSB and BSB transitions get modified as [1],
@@
Unlike the previous case, the Rabi frequency ($\Omega_{0}$) remains unchanged for the carrier transition. Thus, when the control laser is tuned to a sideband transition (BSB or RSB), the effect of the phonon occupancy, $n$, becomes evident in the EIT linewidth through its influence on the Rabi frequency, $\Omega_{c}$. In other words, variations in the EIT linewidth provide an estimate of changes in the ion’s vibrational occupancy. Motivated by this correlation, we define a systematic approach via numerical simulations for determining the ion temperature via phonon occupancy.
@@
In the rotating frame defined by the probe frequency $\omega_{probe}$, the dynamics of the system can be described by the following time-independent Hamiltonian under the electric dipole and rotating-wave approximation (RWA) (with $\hbar$ =1):
@@
$\frac{d\rho}{dt}=$ $-i[H,\rho]+\kappa\left(2a\rho a^{\dagger}-a^{\dagger}a\rho-\rho a^{\dagger}a\right)$ (3)
$+\gamma_{eu}\left(2\sigma_{ue}\rho\sigma_{eu}-\sigma_{ee}\rho-\rho\sigma_{ee}\right)$
$+\gamma_{eg}\left(2\sigma_{ge}\rho\sigma_{eg}-\sigma_{ee}\rho-\rho\sigma_{ee}\right)$
$+\gamma_{b}(n_{\mathrm{th}}+1)\left(2b\rho b^{\dagger}-b^{\dagger}b\rho-\rho b^{\dagger}b\right)$
$+\gamma_{b}n_{\mathrm{th}}\left(2b^{\dagger}\rho b-bb^{\dagger}\rho-\rho bb^{\dagger}\right)$
where $\rho$ represents the density matrix of the ion-cavity system and $\kappa$ denotes the decay rate of the cavity field.
@@
The average phonon number ($\bar{n}$) is related to the temperature of the ion as,
@@
The phonon states are inherent to damping due to fluctuating electric field noise present in real ion trap. To account for this scenario, we considered the trapped ion system such that it is coupled to a thermal bath. Fluctuating electric-field noise induces phonon exchange between the ion and the reservoir, driving the system toward thermal equilibrium [8].
The rate equation for the phonons can be derived as [8],
@@
so close to the ground state ($\bar{n}=0$), $\dot{\bar{n}}(t)=\gamma_{b}n_{th}$. This gives an initial estimation of the heating rate of the ion. Equation 5, also shows that if a long time interval is taken, i.e., at equilibrium, $\bar{n}_{eq}=n_{th}$. So, determining the value of $n_{th}$ allows us to estimate the ion’s temperature owing to Equation 4.
@@ III Analysis of Transmission in Cavity-EIT
To study the system dynamics, we numerically solve the master equation (Equation 3) using the Hamiltonian given in Equation 2 for a single ion strongly coupled to the intracavity field. The probe-field detuning is varied across the cavity resonance while keeping the other system parameters fixed. The steady-state density matrix is obtained for each value of the probe detuning.
@@
From the steady-state solution, the photon population inside the cavity is extracted by evaluating the expectation value of the photon number operator, $\langle a^{\dagger}a\rangle$. The steady-state intra-cavity photon number, which is proportional to the output field leaking through the cavity mirrors, is then used to calculate the transmission spectrum. Figure 2) shows the resulting transmission profile. This spectrum exhibits the characteristic features of previously reported cavity-EIT systems [19, 36, 20, 27].
@@
The cavity linewidth is observed to be narrower than that of an empty cavity, and the vacuum Rabi splitting peaks appear at $\Delta=\pm\sqrt{g^{2}+\Omega_{c}^{2}}$. For all simulations presented in this work, unless otherwise stated, the system parameters are fixed as follows. All quantities are normalized with respect to the cavity decay rate, $\kappa$, which is taken to be $0.4~\mathrm{MHz}$, a typical value for a high-finesse cavity. The coupling-field Rabi frequency, $\Omega_{c}$, is set to $2.5\kappa$, while the atom-cavity coupling strength, $g$, is chosen as $3\kappa$. The decay rates from the excited states, $\gamma_{ge}$ and $\gamma_{ue}$, are taken to be equal to $\kappa$. In addition, motional-state decoherence is included to account for unwanted heating or cooling of the ion arising from phase noise and radiative fluctuations present in realistic ion traps. To model this effect, a weak coupling between the motional state and the environment is introduced, characterized by a coupling constant $\gamma_{b}\approx 0.25\kappa$.
@@ IV Effect of thermal state on cavity EIT linewidth
The usual cavity-EIT formalism discussed and reviewed through multiple theoretical approaches [37, 38, 39] as well as experimental investigations [40, 32] till date has not explicitly considered the ion’s motional state in the realization of the cavity-induced EIT phenomenon. This inherently makes an assumption that the ion is fixed in space and the effect of ion thermalization does not affect the cavity-EIT output. When the ion’s motion is not considered for cavity-based EIT, the Hamiltonian is simplified to,
@@
where the different terms have their usual meaning as described in section II. We refer to this formalism as non-thermal and the Hamiltonian given in Equation 6, the non-thermal cavity-EIT Hamiltonian. The output of the cavity is traced from the average of the photon number operator with the variation of the probe field detuning. The same cavity transmission can also be analytically calculated from the expression[41],
@@
When the motional states are incorporated into the Hamiltonian (Equation 10), it encompasses a more general description of a trapped ion, the cavity-EIT linewidth deviates from the non-thermal case. As the temperature increases, the thermal nature of the ion leads to a higher motional occupancy. This increase in motional occupation modifies the cavity-EIT linewidth by changing $\Omega_{c}$, as discussed in detail in section II.
To compare the cavity-induced EIT linewidths in the thermal (when motional states are included) and non-thermal (when motional states are not considered) cases, we perform simulations for both scenarios using the same values of $\Omega_{c}$ as those employed in the analytical plot in 3(a). The thermal simulations are carried out over a range of temperatures to capture the dependence of the cavity-induced EIT linewidth on the ion temperature. In 3(b), the thermal and non-thermal linewidths are presented, with the analytical linewidths from 3(a) used as the reference on the x-axis.
It is shown that, over the entire range, the non-thermal linewidth remains lower than the thermal linewidth and that the linewidth curves shift toward higher values as the temperature increases. This behavior provides further motivation for quantifying the ion temperature, as it modifies the cavity linewidth, which is highly sensitive to temperature-induced changes in the ion’s motional state.